Shift report 20030511 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun May 11 2003 - 08:01:52 EDT

  • Next message: McBreen Betty: "rfptexp upgrade notice"
    ------------------------------------------------------------
    Date: Sun May 11 08:01:52 2003, Shift: 00:00-08:00
    Supervisor: sjs, with: 
    Log book#14 pages 106 - 106
    
    Shift Summary:
    ------------------------------------------------------------
    Beam for almost 6.5 hours during shift, 
    with only one beam
    dump (voted by STAR and PHENIX).  
    (PHENIX just voted for a dump. 7:53 am).
    Data runs (trig 2): 
    8896(376),
    8897(90),
    8900(312),
    8901(158,still running).
    
    One problem noted during shift: We have lost the
    rate histograms.  These seem to have disappeared 
    sometime last evening.  Could this be related to
    the polarization scaler tests?
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun May 11 2003 - 08:05:39 EDT